Circular No. 6105

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1994ae IN NGC 3370 S. D. Van Dyk and R. R. Treffers, University of California at Berkeley

(UCB); M. W. Richmond, Princeton University; and A. V. Filippenko and Y. B. Paik, UCB, report their discovery of an apparent supernova in NGC 3370 (α = 10^{h}44^{m}23^{s}, δ = +17^{o}32′.3, equinox 1950.0). Observations were conducted as part of the Leuschner Observatory Supernova Search (LOSS), which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. The supernova is located ∼ 30′′.3 west and 6′′.1 north of the galaxy’s nucleus. The discovery image was taken on Nov. 14 UT, and the object is visible at R = 15.4 ± 0.4. A second image taken nearly 2 hr later shows the object at the same position; it is not present to a limiting R magnitude of ∼ 17 in previous LOSS images of the galaxy, the most recent one having been obtained on Nov. 3.

GX 1+4 D. Chakrabarty and T. A. Prince, California Institute of Technology;

M. H. Finger, Universities Space Research Association; and R. B. Wilson, Marshall Space Flight Center, NASA, report for the Compton Gamma Ray Observatory BATSE team: “The accreting x-ray pulsar GX 1+4 has resumed spinning up, reversing the steady spin-down trend that had been observed since 1987. We observed a smooth decrease in the spin-down rate, from (3.8 ± 0.1) × 10^{−12} s^{−2} on June 20 to zero on ∼ Oct. 30. During the interval Nov. 8–11, the spin-up rate was (2.7 ± 0.3) × 10^{−12} s^{−2}, with a mean barycentric pulse frequency of (8.19033 ± 0.00003) × 10^{−3} s^{−1}. The pulsed flux has increased by a factor of three since Oct. 17, when the spin-down rate was (1.5 ± 0.2) × 10^{−12} s^{−2}. On Nov. 10, the 20- to 100-keV phase-averaged pulsed flux was 95 ± 10 mCrab, with a spectrum well fit by an optically-thin thermal bremsstrahlung model F (E) = (A/E)exp(−E/kT), with temperature kT = 35 ± 2 keV and F (50 keV) = (2.6 ± 0.1) × 10^{−4}. The pulsed emission is brighter and harder than observed during a previous outburst (IAUC 5859), and is continuing to increase in intensity.”

COMET MACHHOLZ (1994r) Total visual magnitude estimates by C. S. Morris, Pine Mountain Club,

CA (0.26-m reflector): Oct. 29.32 UT, 10.6; 30.34, 10.5; Nov. 12.47, 10.7.

1994 November 16 Daniel W. E. Green